a freight train leaves the station 2 hours before a passenger train. There are two trains are traveling in the same direction on parallel tracks. If the rate of the passenger train is 11 mph faster than the Freight train, how fast is each train traveling if the passenger train passes the freight train in 10 hours?
Found 2 solutions by mananth, angelchristian16:
Answer by mananth(16949)

a freight train leaves the station 2 hours before a passenger train. There are two trains are traveling in the same direction on parallel tracks. If the rate of the passenger train is 11 mph faster than the Freight train, how fast is each train traveling if the passenger train passes the freight train in 10 hours?
..
freight train = x mph
passenger train = x+11 mph
...
freight train time = 10 +3 = 13 hours
passanger train = 10 hours
..
distance same. because they meet
..
13x=10(x+11)
13x=10x+110
13x-10x=110
3x= 110
/3
x= 36.66 mph speed of freight train
passenger train speed = 36.66 +11 = 47.66 mph

Answer by angelchristian16(1)

Freight Train Data:
time 10+2 hours, rate = r mph, distance = r(12)miles
Passenger Train Data:
time 10 hours, rate = r+11 mph, distance = 10(r+11)miles
Equation:
12r= 10r + 110
2r= 110
R= 55 mph Freight Train Speed
r2+ 11 mph
R2=66 mph Passenger Train Speed